I pharmacists wishes to mix a solution that is 3% minoxidil . She has on hand 70 mL of a 1% solution and wishes to add some 5% solution to obtain the desired 3% solution how much 5% solution should she had
Found 2 solutions by stanbon, fractalier:
Answer by stanbon(75887)

I pharmacists wishes to mix a solution that is 3% minoxidil . She has on hand 70 mL of a 1% solution and wishes to add some 5% solution to obtain the desired 3% solution how much 5% solution should she had
-----
Equation:
mino + mino = mino
0.01*70 + 0.05x = 0.03(70+x)
----
Multiply thru by 100 to get:
70 + 5x = 3*70 + 3x
---
2x = 140
x = 70 mL (amt. of 5% solution needed)
-----
Cheers,
Stan H.
---------

Answer by fractalier(6550)

Call the amount added, x. The set up looks like this
70(.01) + x(.05) = (70+x)(.03)
.7 + .05x = 2.1 + .03x
-1.4 = -.02x
x = 1.4 / .02 = 70 mL
Makes sense too...